# Learning to Generate Vectorized Maps at Intersections with Multiple Roadside Cameras

### v1

@@ VI-A Real-world Dataset
In this study, we developed a custom dataset to facilitate our research objectives and named Navinfo dataset, as depicted in Fig. II. The dataset is primarily comprised of video surveillance data from traffic intersections. Each intersection is monitored by cameras positioned to capture images from four distinct orientations. Each camera boasts a resolution of 2 million pixels, allowing for high-quality imagery. The combined view from these four cameras provides comprehensive coverage of the entire intersection as well as a portion of the roadways. In addition to image information, Navinfo dataset also encompasses several important information. Firstly, the camera metadata includes intrinsic parameters, quaternions, Euler angles (pitch, yaw, roll), geographical coordinates (latitude and longitude of WGS84 coordinate system [35]), elevation, and real-world coordinates relative to the camera’s position. Secondly, the map information includes detailed descriptions of ground elements such as lane markings, road edges, pedestrian crossings, lane dividers, road boundaries, directional arrows, inscriptions on the road surface, ground signs, and traffic islands. Additionally, elevation elements such as traffic lights, signage, and guardrails are included, all with precise geographic coordinates and shape information. By applying camera parameters to process the images, we establish a one-to-one correspondence with the map information. In terms of Navinfo dataset scale, data collection was conducted across 4,000 intersections in four major cities: 1,246 intersections in Beijing, 1,121 in Shanghai, 988 in Guangzhou, and 645 in Shenzhen. A total of 16,000 roadside camera datasets were compiled, which include 160,000 one-minute video clips captured periodically every 25 hours per camera, all of which were recorded during daytime conditions. This dataset enables the generation of 400,000 image groups, amounting to a total of 1.6 million images, alongside 16,000 groups of ground truth data. Importantly, the map information (ground truth) was meticulously collected by NavInfo using a laser-equipped vehicle, thereby ensuring a high degree of precision and reliability in the dataset.
@@ VI-C Edge Deployment
This project seeks to deploy a comprehensive set of algorithms on edge devices, with the Jetson AGX Orin selected as the target platform. The Jetson AGX Orin provides exceptional computational capabilities, boasting up to 275 TOPS of AI performance in INT8 precision—an impressive eightfold performance increase compared to its predecessor, the Jetson AGX Xavier. For video processing, the Jetson AGX Orin supports video encoding for up to 16 channels of 2-megapixel video at 30 FPS, as well as video decoding for up to 24 channels of 2-megapixel video at 30 FPS. Its power consumption ranges from 15W to 60W, effectively accommodating the real-time operational demands of the algorithms presented in this study. The Jetson AGX Orin is compatible with the NVIDIA JetPack SDK. This infrastructure provides a robust foundation for the deployment of our algorithms. The initial step involved configuring the JetPack SDK to enable the utilization of CUDA, cuDNN, and TensorRT, alongside setting up the GCC compilation environment. Subsequently, the algorithm model was converted to a TensorRT-compatible format. Specifically, the initial conversion process involved transforming the model from PyTorch (.pth) format to ONNX format, necessitating support from the MMDeploy Toolbox and the implementation of custom nodes. Following this step, the ONNX model was converted to a TensorRT-compatible format utilizing the trtexec script provided by TensorRT. To optimize the algorithm’s efficiency without compromising its accuracy, the inference model was converted to FP16 precision. The final implementation included data reading, model preprocessing, and postprocessing conducted in C++, culminating in the rendering of visual outcomes based on the results obtained from the postprocessing procedures. Upon deployment, The MRC-VMap, which uses ResNet50 as its backbone network, occupies approximately 1.6GB of GPU memory and achieves an inference speed of 20 FPS, while maintaining an overall accuracy degradation of less than 1%.
@@ VI-D Comparison Results
This research represents the initial effort to investigate the generation of high-definition maps utilizing multi-camera data from intersections, leading to a scarcity of widely recognized comparative methods in the existing literature. However, given that the data structure of the Navinfo dataset is modeled after the well-established NuScenes dataset format, we are able to conduct comparative experiments on the Navinfo dataset by assessing several classic open-source algorithms, such as HDMapNet. This comparison follows the algorithm performance evaluation approach employed in studies utilizing the NuScenes dataset. Table V-C summarizes the experimental results from various algorithms, including HDMapNet, VectorMapNet, MapTR, MRC-VMap, and VI-Map, on the Navinfo dataset. The comparison primarily focuses on the accuracy of three key elements: pedestrian crossings, lane dividers, and road boundaries. The results indicate that the MRC-VMap significantly improves the quality of high-definition map generation, achieving an increase of over 3 mAP in pure visual mode. This outcome underscores the effectiveness of the MRC-VMap algorithm in generating high-quality, high-definition maps within the framework of the Navinfo dataset. Furthermore, we assessed the efficiency of MRC-VMap, finding it to be 28% more efficient than the current leading algorithm, MapTR, when executed on an RTX 3090 GPU. Lastly, we implemented the VI-Map method on the Navinfo dataset, utilizing both camera and LiDAR point cloud data as inputs; its accuracy was reported to be merely 1.3 mAP higher than that of MRC-VMap. The prediction results of the vectorized high-definition maps are illustrated in Fig. VI-A.
@@ VI-E1 Ablation on FPN Variants
Performance comparison of different backbones and BEV acquisition methods.
ResNet18 GKT 49.1 26.1 LSS 48.3 25.6 BEV-Learner 49.0 40.9 ResNet50 GKT 63.2 13.8 LSS 63.1 13.4 BEV-Learner 63.0 18.2
@@ VI-E2 Ablation on BEV Extractor
Evaluating various methodologies for transforming perspective view image features into the BEV space features, Table VI-E1 presents the results for methods including LSS, GKT, and the proposed BEV-Learner approach. The experimental outcomes demonstrate that the BEV-Learner method, regardless of whether ResNet18 or ResNet50 is used, consistently exhibits higher operational efficiency while achieving accuracy levels comparable to the current best-performing GKT method. Specifically, the implementation of the BEV-Learner method led to a 57% improvement in efficiency when utilizing ResNet18, and a 32% enhancement when using ResNet50, all while maintaining accuracy on par with the GKT approach.
: TABLE IPerformance comparison of different quantities of inputs.
@@ VI-E3 Ablation Study on Single-Camera Mode
Table LABEL:tab:4 exhibits the performance of MRC-VMap when utilizing data from a single camera as input in an intersection scenario. The analysis reveals that when ResNet50 serves as the backbone of the algorithm, the single-camera mode presents certain advantages in terms of efficiency when compared to the four-camera mode (as detailed in Table V-C). However, it is noteworthy that there remains a substantial discrepancy in accuracy, with a reduction of 15.1 mAP evident in the single-camera configuration.

### v3

@@ V-C Training Loss Functions
VectorMapNet C ResNet50 110 6 36.1 47.3 39.3 40.9 2.9 MapTR (nano) C ResNet18 110 6 39.6 49.9 48.2 45.9 25.1 MapTR (tiny) C ResNet50 24 6 46.3 51.5 53.1 50.3 11.2 MapTR (tiny) C ResNet50 110 6 56.2 59.8 60.1 58.7 11.2 MRC-VMap (nano) C ResNet18 110 6 39.8 50.4 48.4 46.2 38.1 MRC-VMap (tiny) C ResNet50 24 6 46.2 51.7 53.6 50.5 14.7 MRC-VMap (tiny) C ResNet50 110 6 58.1 62.5 61.2 60.6 14.7
  • Table LABEL:tab:1 presents a comparison of our method with state-of-the-art approaches ( [[\@@bibref{}{liu2023vectormapnet}{}{}], cite],  [[\@@bibref{}{maptr}{}{}], cite]) utilizing the NuScenes validation set. The definitions of the fields in this table are identical to those in Table  V-C. The quantitative findings demonstrate that MRC-VMap (tiny) significantly outperforms the other methods listed in the table. MRC-VMap (nano) achieves state-of-the-art performance, running at an impressive 38.1 FPS.
@@ VI-A Real-world Dataset
In this study, we developed a custom dataset to facilitate our research objectives and named Navinfo dataset, as depicted in Fig. II. The dataset is primarily comprised of video surveillance data from traffic intersections. Each intersection is monitored by cameras positioned to capture images from four distinct orientations. Each camera boasts a resolution of 2 million pixels, allowing for high-quality imagery. The combined view from these four cameras provides comprehensive coverage of the entire intersection as well as a portion of the roadways. In addition to image information, Navinfo dataset also encompasses several important information. Firstly, the camera metadata includes intrinsic parameters, quaternions, Euler angles (pitch, yaw, roll – these are labeled once and may vary from actual values due to the long data collection period), geographical coordinates (latitude and longitude of WGS84 coordinate system [35]), elevation, and real-world coordinates relative to the camera’s position. Secondly, the map information includes detailed descriptions of ground elements such as lane markings, road edges, pedestrian crossings, lane dividers, road boundaries, directional arrows, inscriptions on the road surface, ground signs, and traffic islands. Additionally, elevation elements such as traffic lights, signage, and guardrails are included, all with precise geographic coordinates and shape information. By applying camera parameters to process the images, we establish a one-to-one correspondence with the map information. In terms of Navinfo dataset scale, data collection was conducted across 4,000 intersections in four major cities: 1,246 intersections in Beijing, 1,121 in Shanghai, 988 in Guangzhou, and 645 in Shenzhen. A total of 16,000 roadside camera datasets were compiled, which include 160,000 one-minute video clips captured periodically every 25 hours per camera, all of which were recorded during daytime conditions. This dataset enables the generation of 400,000 image groups, amounting to a total of 1.6 million images, alongside 16,000 groups of ground truth data. Importantly, the map information (ground truth) was meticulously collected by NavInfo using a laser-equipped vehicle, thereby ensuring a high degree of precision and reliability in the dataset.
@@ VI-C Edge Deployment
This project aims to implement a range of algorithms on edge devices, selecting the Jetson AGX Orin as the target platform. It is compatible with NVIDIA JetPack SDK. The process begins with configuring the JetPack SDK to access CUDA, cuDNN, and TensorRT, and setting up the GCC compilation environment. The algorithm model was converted to a TensorRT-compatible format and optimized to FP16 precision to maintain accuracy while improving efficiency. Finally, the implementation involved data reading, model preprocessing, and postprocessing in C++, with visual results rendered based on the postprocessing outcomes. Upon deployment, The MRC-VMap, which uses ResNet50 as its backbone network, occupies approximately 1.6GB of GPU memory and achieves an inference speed of 20 FPS, while maintaining an overall accuracy degradation of less than 1%.
@@ VI-D Comparison Results
This research represents the initial effort to investigate the generation of high-definition maps utilizing multi-camera data from intersections, leading to a scarcity of widely recognized comparative methods in the existing literature. However, given that the data structure of the Navinfo dataset is modeled after the well-established NuScenes dataset format, we are able to conduct comparative experiments on the Navinfo dataset by assessing several classic open-source algorithms, such as HDMapNet. This comparison follows the algorithm performance evaluation approach employed in studies utilizing the NuScenes dataset. Table V-C summarizes the experimental results from various algorithms, including HDMapNet, VectorMapNet, MapTR, MRC-VMap, and VI-Map, on the Navinfo dataset. The comparison primarily focuses on the accuracy of three key elements: pedestrian crossings, lane dividers, and road boundaries. The results indicate that the MRC-VMap significantly improves the quality of high-definition map generation, achieving an increase of over 3 mAP in pure visual mode. This outcome underscores the effectiveness of the MRC-VMap algorithm in generating high-quality, high-definition maps within the framework of the Navinfo dataset. Furthermore, we assessed the efficiency of MRC-VMap, finding it to be 28% more efficient than the current leading algorithm, MapTR, when executed on an RTX 3090 GPU. Lastly, we implemented the VI-Map method on the Navinfo dataset, utilizing both camera and LiDAR point cloud data as inputs; its accuracy was reported to be merely 1.3 mAP higher than that of MRC-VMap. The prediction results of the vectorized high-definition maps are illustrated in Fig. VI-A. Table LABEL:tab:1 summarizes the experimental results from various algorithms, including VectorMapNet, MapTR and MRC-VMap on the NuScenes dataset. The results show that MRC-VMap achieves a 1.9 mAP increase and its operational efficiency has improved by 31% compared to MapTR. This also indicates that MRC-VMap can enhance the quality and efficiency of high-definition map generation.
@@ VI-E1 Ablation on FPN Variants
: TABLE IPerformance comparison of different backbones and BEV acquisition methods.
Backbone Method mAP FPS
ResNet18 GKT 49.1 26.1
LSS 48.3 25.6
BEV-Learner 49.0 40.9
ResNet50 GKT 63.2 13.8
LSS 63.1 13.4
BEV-Learner 63.0 18.2
@@ VI-E2 Ablation on BEV Extractor
Evaluating various methodologies for transforming perspective view image features into the BEV space features, Table LABEL:tab:3 presents the results for methods including LSS, GKT, and the proposed BEV-Learner approach. The experimental outcomes demonstrate that the BEV-Learner method, regardless of whether ResNet18 or ResNet50 is used, consistently exhibits higher operational efficiency while achieving accuracy levels comparable to the current best-performing GKT method. Specifically, the implementation of the BEV-Learner method led to a 57% improvement in efficiency when utilizing ResNet18, and a 32% enhancement when using ResNet50, all while maintaining accuracy on par with the GKT approach.
: TABLE IIPerformance comparison of different quantities of inputs.
@@ VI-E3 Ablation Study on Single-Camera Mode
Table LABEL:tab:4 exhibits the performance of MRC-VMap when utilizing data from a single camera as input in an intersection scenario. The analysis reveals that when ResNet50 serves as the backbone of the algorithm, the single-camera mode presents certain advantages in terms of efficiency when compared to the four-camera mode (as detailed in Table LABEL:tab:1). However, it is noteworthy that there remains a substantial discrepancy in accuracy, with a reduction of 15.1 mAP evident in the single-camera configuration.